(a) The height of trees a normally distributed with mean 160 cm and standard deviation is 8 cm.
Find the probability data randomly selected tree has a height greeter than 170 cm
Find the probability that tree has a height greeter than 180 cm.
Give intuitive reason why your answers are relative.
(b) Find median and mode.
1
Expert's answer
2020-08-05T18:34:38-0400

Solution a)


"\mu = 160"

"\sigma = 8"


Let tree height be X


"p(X > 170)"


"Z = \frac{x-\mu}{\sigma}"

"= \frac{170-160}{8} = 1.25"


At Z = 1.25 the probability p(X>170) = 0.1056


Answer: 0.1056



"p(X>180)"


"Z = \frac{x-\mu}{\sigma}"

"Z=\frac{180−160}{8} = 2.5"

At Z = 2.5 the probability p(X>180) = 0.0062


Answer: 0.0062


These solutions are relative since the calculated probabilities are based on the unique distribution of the data defined by the mean (160) and standard deviation (8)


Solution b)


For a normal curve


"mean = median = mode"


Answer: Median = Mode = 160